what is amount of liabilities? net inc. 50000, return on assets 12.12%, return on equit 40%
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Total assets = net income / return on assets = 50000/12.12% = $412500 Equity = net income / return on equity = 50000/4